OpenStudy (anonymous):

For her exercise today, Mary plans to both run and swim. Let r be the number of laps she runs and let be the number of laps she swims. Each lap she runs takes her 4 minutes, and each lap she swims takes her 3 minutes. She wants to exercise for at least 30 minutes today. Using the values and variables given, write an inequality describing this.

OpenStudy (anonymous):

*1 Let r be the number of laps she runs. *2 Let j be the number of laps she swims. Ok? Then \[\frac{4}{60} + \frac{ 3 }{ 60 } \ge \frac{ 30 }{ 60 }\]